OpenStudy (anonymous):

Amy leaves a rest stop at 2 PM traveling west, at 45 miles per hour. Sarah leaves 2 hours later, also traveling west, but at 55 miles per hour. At what time will Amy catch up with Sarah? You solve the problem by setting up the table and writing the equation 55(t - 2)= 45t. When you solve, you get 11. What is the answer to the problem?

OpenStudy (anonymous):

miles for amy: 45 * h + 45*2<- 2hours advantage miles for Sarah: 55 * h So solve the following for h: 45 * h + 45*2 = 55 * h h=18 So it would take 18 hours to catch up